A condensed balance sheet for Bradford Corporation prepared at the end of the year appears as follows.

Assets Liabilities & Stockholders' Equity
Cash $ 104,000 Notes payable (due in 6 months) $ 45,000
Accounts receivable 100,000 Accounts payable 122,000
Inventory 270,000 Long-term liabilities 321,000
Prepaid expenses 60,000 Capital stock, $5 par 300,000
Plant & equipment (net) 570,000 Retained earnings 406,000
Other assets 90,000
Total $ 1,194,000 Total $ 1,194,000

During the year, the company earned a gross profit of $1,116,000 on sales of $2,950,000. Accounts receivable, inventory, and plant assets remained almost constant in amount throughout the year, so year-end figures may be used rather than averages.

a. Compute the current ratio. (Round your answer to 2 decimal place.)

b. Compute the quick ratio. (Round your answer to 2 decimal place.)

c. Compute the working capital.

d. Compute the debt ratio. (Round your percentage answers to nearest whole percent. i.e. 0.1234 as 12%.)

e. Compute the accounts receivable turnover (all sales were on credit). (Round your answer to 2 decimal places.)

f. Compute the inventory turnover. (Round your answer to 2 decimal places.)

g. Compute the book value per share of capital stock. (Round your answer to 2 decimal places.)
Answer #1
a.
Cash 104000
Accounts receivable 100000
Inventory 270000
Prepaid expenses 60000
Current assets 534000
Notes payable (due in 6 months) 45000
Accounts payable 122000
Current liabilities 167000
Current ratio = Current assets / Current liabilities = 534000 / 167000 3.20
b.
Quick ratio = ( Current assets - Inventory - Prepaid expenses ) / Current liabilities = ( 534000 - 270000 - 60000 ) / 167000 1.22
c.
Current assets 534000
(-) Current liabilities 167000
Working capital 367000
d.
Current liabilities 167000
(+) Long-term liabilities 321000
Total liabilities 488000
Debt ratio = Total liabilities / Total assets = 488000 / 1194000 41%
e.
Accounts receivables turnover = Credit sales / Accounts receivable = 2950000 / 100000 29.50
f.
Cost of goods sold = Sales - Gross profit = 2950000 - 1116000 1834000
Inventory turnover = Cost of goods sold / Inventory = 1834000 / 270000 6.79
g.
Capital stock, $5 par 300000
(+) Retained earnings 406000
Total stockholders' equity 706000
Shares outstanding = Capital stock / Par value per share = 300000 / 5 60000
Book value per share of capital stock = Total stockholders' equity / Shares outstanding = 706000 / 60000 11.77